OpenStudy (anonymous):

Would anyone be willing to help me with a pre-cal question? I'll post it below

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} -4(1/3)^n-1\]

OpenStudy (anonymous):

a. Write the first four terms of the series b. Does the series diverge or converge? c. If the series has a sum, find the sum

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hi there :) Is the negative 1 supposed to be in the exponent? \[\Large\rm \sum_{n=1}^{\infty}-4\left(\frac{1}{3}\right)^{n-1}\]Like this?